OpenStudy (anonymous):

9c+4=3c-8 i need to figure out the value of c . step by step .

OpenStudy (anonymous):

9c-3c=-8-4 6c = -12 c=-2

OpenStudy (anonymous):

6c=-12===>c =-2

OpenStudy (anonymous):

how do you get -8 ?